if two straight line represented by the equation x2(tan2 + cos2 )-2xytan +y2 sin2=0 makes angle A and B with x axis respectively ,than show that tan A - tan B


1
Expert's answer
2021-09-02T07:25:24-0400
"x ^2(\tan^2θ+\cos^2θ)−2xy\tanθ+y^2\sin^2θ=0"

"y^2\sin^2θ−2xy\tanθ+x ^2(\tan^2θ+\cos^2θ)=0"

"y=\dfrac{-x\tanθ\pm\sqrt{(x\tanθ)^2-\sin^2θ(x ^2(\tan^2θ+\cos^2θ))}}{\sin^2θ}"

"=\dfrac{-x\tanθ\pm x\cos \theta\sqrt{\tan^2θ-\sin^2θ}}{\sin^2θ}"

"=\dfrac{-x\tanθ\pm x\sin^2 \theta}{\sin^2θ}"

"y_1=(\dfrac{-\tan\theta}{\sin^2θ}+1)x=(\tan A)x"

"y_2=(\dfrac{-\tan\theta}{\sin^2θ}-1)x=(\tan B)x"

Then


"\tan A-\tan B=(\dfrac{-\tan\theta}{\sin^2θ}+1)-(\dfrac{-\tan\theta}{\sin^2θ}-1)"

"\tan A-\tan B=2"